OpenStudy (anonymous):

A triangular prism has a volume of 10 cm3. If each dimension of the prism is multiplied by four, what is the new volume of the prism? 320 cm3 80 cm3 160 cm3 640 cm3

OpenStudy (anonymous):

give me a sec to solve

OpenStudy (anonymous):

thank u im redoing them to bost my grade

OpenStudy (anonymous):

answer is 640 cm^3 how??? because solution: V= 1/2* l *w*h given V = 10 cm^3 but here given , new vol= 1/2* 4l*4w*4h =64 * 1/2* l*w*h =64*V =64*10 =640 cm^3

OpenStudy (anonymous):

wow ur good can u help me again??

OpenStudy (anonymous):

sure i'll try

OpenStudy (anonymous):

Each dimension of a rectangular prism is increased by 30%. What percent of the original volume is the volume of the new prism? 9 percent 169 percent 2.7 percent 219.7 percent

OpenStudy (anonymous):

ok wait lemme solve

OpenStudy (anonymous):

answer is 219.7% how lets prove....

OpenStudy (anonymous):

l*w*h = V if we use h= 2 w= 3 l= 4 V= 24 units^3 increase by 30% we get, h= 2.6 w= 3.9 l= 5.2 v =52.73 units^3 which is 219.7% time of the original volume original volume * 219.7% => 24 * 219.7% = 52.728 hence proved